Three point charges are located at the corners of an equilateral triangle, whose side l = 0.5 m. The charges have magnitude -7.00 μC, 2.00 μC and 2.00 μC respectively. (a) Calculate the total electric potential energy of these charges. (b) How much work must be done to move the 2.00 μC charge to infinity, leaving the other two charges in place? (d) Find the (net) electric field at the midpoint between the 2.00 μC and -2.00 μC charges.
Answer #1

2, =-7li y Tofal electric potential energy - 2 MC =q3 azale o-son.al L VE Ac @ ka, qz kq, 93 Kq263 r t r = 9.99x109 x ()x5+1-Emet - Senter = 5(575360)2 + (335636-6667) N/c NIC こ 6660963911 = 666 096 · 38